Weather: Same as yesterday. Haze, sunny, calm winds. High in the low 40’s. NWS shows a large blocking high pressure system centered over the eastern pacific, just west of Washington which is affecting the entire Western US. Looks like we’ll be under this system for the rest of the week.
Drove Matt to the airport while John attended a telecon this morning. I moderated Holger’s EOL seminar in the afternoon while John launched the 415p radiosonde. Thanks John!
- Nothing disruptive to report in the network connections - Nagios has been green all day for all ISFS sites. No spikes to filter, nothing odd in the data itself to report.
- Steve asked us to find SHT sensors in the base trailer. We *think* found 13 of them in a pink ziploc (please confirm):
- LC - still intermittent sonic data concurrent with the IRGA starting ~ Jan 28 (See Matt's blog post). We plan to replace the EC100 tomorrow and also take soil sample #3 while there. Apparently, there is a 30% chance of this working.
- If there is a sequence of procedures to follow to replace the EC100, please let us know. Otherwise, John's recollection is it's a plug and play.
- Speaking of soil samples, I found that the door to the oven that dries the samples opened mysteriously overnight. Since we’re simply drying out the samples, this incident shouldn’t affect the measurements. However, I got paranoid and when I checked it 10 minutes later it was open again! I placed a barrier to keep the door shut. I’ll be doing this from now on.
IOP 5 is planned for tomorrow, noon Wednesday to noon Thursday. Classified as "Moisture surge".